OpenStudy (anonymous):

What is the value of x? I needs help. c: fan and medal for the answer WITH an explanation. c: thank you

OpenStudy (anonymous):

example task?

OpenStudy (anonymous):

\[\frac{ x }{ 5} +6=36\]

OpenStudy (anonymous):

Its kinda like this, I used different numbers. Tell me if the numbers dont work with the equation, because I can change them if they dont...

OpenStudy (here_to_help15):

ok so write 6 as a fraction

OpenStudy (here_to_help15):

6/1

OpenStudy (here_to_help15):

and if im correct you flip x/5

OpenStudy (here_to_help15):

nvm hes correct

OpenStudy (here_to_help15):

i was thinking of something else more complicated

OpenStudy (zephyr141):

you end up with x/5=30 so you divide 30 by 5 to get x which is just 6.

OpenStudy (anonymous):

Thank you guys. :) I will fan you later, because my internet is being sucky. And okay, @zephyr141

OpenStudy (anonymous):

So the answer to the equation is just 6?

OpenStudy (zephyr141):

yeah. i messed up the math in the first part.

OpenStudy (anonymous):

lol, okay... thank you!

OpenStudy (zephyr141):

ugh!!!! it's too early. i'm doing it wrong again.

OpenStudy (zephyr141):

x=150. (x/5)+6=36 subtract 6 from both sides (x/5)=30 now multiply both sides by 5 to get x=150. sorry for the wrong answer before